OpenStudy (anonymous):

Solve P = 2w + 2h for w.

OpenStudy (calculusxy):

You are trying to find the perimeter.

OpenStudy (calculusxy):

w=P1/2-h

OpenStudy (johnweldon1993):

So since we can't actually solve for 'w' because we have too many unknowns...what we would do...is isolate 'w'....w = ..... So P = 2w + 2h Subtract 2h from both sides... P - 2h = 2w Now divide both sides by 2....because this will leave just 'w' \[w = \frac{ 1 }{ 2 }P - h\]